# Training and Evaluating with Human Label Variation: An Empirical Study

### v4

@@ Abstract
Human label variation (HLV) challenges the standard assumption that a labelled instance has a single ground truth, instead embracing the natural variation in human annotation to train and evaluate models. While various training methods and metrics for HLV have been proposed, it is still unclear which methods and metrics perform best in what settings. We propose new evaluation metrics for HLV leveraging fuzzy set theory. Since these new proposed metrics are differentiable, we then in turn experiment with employing these metrics as training objectives. We conduct an extensive study over 6 HLV datasets testing 14 training methods and 6 evaluation metrics. We find that training on either disaggregated annotations or soft labels performs best across metrics, outperforming training using the proposed training objectives with differentiable metrics. We also show that our proposed soft micro F1 score is one of the best metrics for HLV data. [^1] Code is available at: https://github.com/kmkurn/train-eval-hlv
@@ 1 Introduction
We find that two of the simplest methods for HLV training perform surprisingly well, outperforming more complex methods including training with each soft metric as the objective and attaining the best performance in most cases. The first method considers each annotation of an instance as a separate instance-label pair. Training is then performed on this disaggregated data where the same instance may occur multiple times with a different label each. The second method trains the model to predict the full label distribution induced by the instance’s annotations using the standard cross-entropy loss. This is in contrast to the standard model training where the prediction target is the mode of the label distribution.
@@ 3 Evaluation Metrics for HLV
Let $P_{ik}\geq 0$ denote the true human judgement for example $i$ and class $k$, i.e. proportion of humans labelling example $i$ with class $k$. Let $Q_{ik}\geq 0$ denote the value of $P_{ik}$ predicted by a model. An evaluation metric $m$ is a function such that the scalar $m(\mathbf{P},\mathbf{Q})$ measures how well (i.e., a positive orientation) $\mathbf{Q}$ captures the judgement distribution in $\mathbf{P}$. To complete our notation, let $N$ and $K$ denote the number of examples and classes respectively.
@@ 6.1 Experimental Setup
To create the ranking of training methods produced by human judgements, given a language model (RoBERTa or LLaMA) we randomly sample a pair of methods (e.g. ReL and SL) and present the instance text (i.e. a legal problem) and two results (i.e. distribution of the areas of law) [^19] Note that because it is multilabel task, the distribution doesn’t sum to one. We only show areas of law whose probability exceeds a threshold of $0.1$. We choose this threshold because it results in similar numbers of areas of law whose probability exceeds the value across all methods. produced by the two methods, and then ask the lawyers which result is more accurate. The lawyers can choose either one of them, both, or neither of them. See Appendix Figure˜7 for an illustration of the annotation task. In total, we collect 1.2K pairwise judgements for both RoBERTa and LLaMA. On average, each instance is annotated by 2.3 lawyers and each pair of methods has 8 judgements.
To make sure the lawyers form their own expectation on the areas of law relevant to an instance before making this pairwise selection, we first show the instance text and ask them to select the most relevant areas of law (Figure˜6 in the Appendix). After completing this task, we then ask them to make the pairwise selection. We conduct several pilot studies with the lawyers to refine the interface design and ensure the task is clear to them.
@@ 6.2 Results
Figure˜5 shows that not all metrics exhibit a strong linear relationship with human judgements. For example, both hard F1 scores and entropy correlation show almost no relationship for LLaMA. All metrics except entropy correlation show a strong linear relationship for RoBERTa. This suggests that entropy correlation is a poor HLV metric.
Table˜5 confirms that all metrics are positively correlated with human judgements but with different magnitudes. As before, both hard F1 scores and entropy correlation are very weakly correlated with human judgements for LLaMA, and entropy correlation is the weakest for RoBERTa. Furthermore, it is the only metric whose correlation is not statistically significant. Across both pretrained models, both PO-JSD and soft micro F1 consistently have the strongest correlation, but each is best for different models: PO-JSD is best for LLaMA while soft micro F1 score is best for RoBERTa. This finding suggests that both metrics are good for evaluation in the HLV context.
@@ Acknowledgements.
This research is supported by the Australian Research Council Linkage Project LP210200917222222https://dataportal.arc.gov.au/NCGP/Web/Grant/Grant/LP210200917 and funded by the Australian Government. This research is done in collaboration with Justice Connect, an Australian public benevolent institution. [^23] As defined by the Australian government: https://www.acnc.gov.au/charity/charities/4a24f21a-38af-e811-a95e-000d3ad24c60/profile We thank Kate Fazio, Tom O’Doherty, and Rose Hyland from Justice Connect for their support throughout the project. We thank David McNamara, Yashik Anand, and Mark Pendergast also from Justice Connect for the development of the meta-evaluation annotation interface. This research is supported by The University of Melbourne’s Research Computing Services and the Petascale Campus Initiative.

### v5

@@ Abstract
Human label variation (HLV) challenges the standard assumption that a labelled instance has a single ground truth, instead embracing the natural variation in human annotation to train and evaluate models. While various training methods and metrics for HLV have been proposed, it is still unclear which methods and metrics perform best in what settings. We propose new evaluation metrics for HLV leveraging fuzzy set theory. Since these new proposed metrics are differentiable, we then in turn experiment with employing these metrics as training objectives. We conduct an extensive study over 6 HLV datasets testing 14 training methods and 6 evaluation metrics. We find that training on either disaggregated annotations or soft labels performs best across metrics, outperforming training using the proposed training objectives with differentiable metrics. We also show that our proposed soft micro F1 score is one of the best metrics for HLV data. [^1] Code and supplementary materials are available at: https://github.com/kmkurn/train-eval-hlv
@@ 1 Introduction
We find that two of the simplest methods for HLV training perform surprisingly well, outperforming more complex methods including training with each soft metric as the objective and attaining the best performance in most cases. The first method considers each annotation of an instance as a separate instance–label pair. Training is then performed on this disaggregated data where the same instance may occur multiple times with a different label each. The second method trains the model to predict the full label distribution induced by the instance’s annotations using the standard cross-entropy loss. This is in contrast to the standard model training where the prediction target is the mode of the label distribution.
@@ 3 Evaluation Metrics for HLV
Let $P_{ik}\geq 0$ denote the human judgement for example $i$ and class $k$, i.e. proportion of humans labelling example $i$ with class $k$. Let $Q_{ik}\geq 0$ denote the value of $P_{ik}$ predicted by a model. An evaluation metric $m$ is a function such that the scalar $m(\mathbf{P},\mathbf{Q})$ measures how well (i.e., a positive orientation) $\mathbf{Q}$ captures the judgement distribution in $\mathbf{P}$. To complete our notation, let $N$ and $K$ denote the number of examples and classes respectively.
@@ 6.1 Experimental Setup
[…]
To make sure the lawyers form their own expectation on the areas of law relevant to an instance before making this pairwise selection, we first show the instance text and ask them to select the most relevant areas of law (Figure˜6 in the Appendix). After completing this task for an instance text, we then ask them to make the pairwise selection. The two tasks must be completed for an instance text before the lawyers can move to the next instance text. We conduct several pilot studies with the lawyers to refine the interface design and ensure the task is clear to them.
@@ 6.2 Results
We draw the regression plots between method performance as given by the evaluation metrics and method scores produced by human judgements in Figure˜5. The figure shows that not all metrics exhibit a strong linear relationship with human judgements. For example, both hard F1 scores and entropy correlation show almost no relationship for LLaMA. All metrics except entropy correlation show a strong linear relationship for RoBERTa. This suggests that entropy correlation is a poor HLV metric.
In Table˜5, we show the Pearson correlation coefficients for all metrics along with the $p$-values. The table confirms that all metrics are positively correlated with human judgements but with different magnitudes. As before, both hard F1 scores and entropy correlation are very weakly correlated with human judgements for LLaMA, and entropy correlation is the weakest for RoBERTa. Furthermore, it is the only metric whose correlation is not statistically significant. Across both pretrained models, both PO-JSD and soft micro F1 consistently have the strongest correlation, but each is best for different models: PO-JSD is best for LLaMA while soft micro F1 score is best for RoBERTa. This finding suggests that both metrics are good for evaluation in the HLV context.
@@ Acknowledgements.
This research is supported by the Australian Research Council Linkage Project LP210200917232323https://dataportal.arc.gov.au/NCGP/Web/Grant/Grant/LP210200917 and funded by the Australian Government. This research is done in collaboration with Justice Connect, an Australian public benevolent institution. [^24] As defined by the Australian government: https://www.acnc.gov.au/charity/charities/4a24f21a-38af-e811-a95e-000d3ad24c60/profile We thank Kate Fazio, Tom O’Doherty, and Rose Hyland from Justice Connect for their support throughout the project. We thank David McNamara, Yashik Anand, and Mark Pendergast also from Justice Connect for the development of the meta-evaluation annotation interface. This research is supported by The University of Melbourne’s Research Computing Services and the Petascale Campus Initiative.